can someone help me on question 1? Because I got incorrect answer

Required information

[The following information applies to the questions displayed below.]

Federated Manufacturing Inc. (FMI) produces electronic components in three divisions: industrial, commercial, and consumer products. The commercial products division annually purchases 10,000 units of part 23–6711, which the industrial division produces for use in manufacturing one of its own products. The commercial division is growing rapidly. The commercial division is expanding its production and now wants to increase its purchases of part 23–6711 to 15,000 units per year. The problem is that the industrial division is at full capacity. No new investment in the industrial division has been made for some years because top management sees little future growth in its products, so its capacity is unlikely to increase soon.

The commercial division can buy part 23–6711 from Advanced Micro Inc. or from Admiral Electric, a customer of the industrial division now purchasing 650 units of part 88–461. The industrial division's sales to Admiral would not be affected by the commercial division’s decision regarding part 23–6711.

Industrial Division:
Data on part 23–6711:
Price to commercial division $ 185
Variable manufacturing costs 155
Price to outside buyers 205
Data on part 88–461:
Variable manufacturing costs $ 65
Sales price 95
Other Suppliers of Part 23–6711:
Advance Micro Inc., price $ 200
Admiral Electric, price 210

Required:

1. What is the proper decision regarding where the commercial division should purchase the additional 5,000 parts, and what is the correct transfer price?

2. Assume that the industrial division’s sales to Admiral will be canceled if the commercial division does not buy from Admiral. What would be the unit cost to FMI in this case, and would the desired transfer price change?

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2

What is the proper decision regarding where the commercial division should purchase the additional 5,000 parts, and what is the correct transfer price?

Cost to FMI
Buy from Advanced Microselected answer correct $1,025,000 per unit
Transfer pricenot attempted $205 per unit
Answer #1

1) OPTION 1 :

When the supplying division operates at full capacity, the minimum transfer price would be:
Marginal Cost per unit + opportunity cost per unit

Since the supplying division is operating at full capacity, it has no incentive to sell the goods to the purchasing division at a price lower than the market price. If the internal order is accepted, capacity is diverted towards the sale. Hence the supplying division would additionally charge the lost contribution from external sales that had to be curtailed.

As the industrial division of FMI is currently supplying 10,000 units of part 23-6711 and it is operating at its full capacity and it is also not given that how much is the capacity of industrial division to produce PART 23-6711 so it's assumed that atleast the industrial division has the capacity of producing 15,000 units of part 23-6711.

In that case minimum transfer price that should be charged by industrial division is

Marginal cost per unit = 155
Add: Contribution to be lost from reduced external sales = 50 (205-155)
Minimum transfer price = $205

Total cost of additional 5,000 units = 5000 units * 205 / unit = $10,25,000

OPTION 2

Purchase from Advance micro Inc. = $200 per unit

Total cost of additional 5,000 units = 5000 units * 200/ unit = $10,00,000

OPTION 3

Purchase from Admiral electric = $210 per unit.

Total cost of additional 5,000 units = 5000 units * 210 / unit = $10,50,000

After analyzing the above three options it is advisable to purchase additional 5,000 units of part 23-6711 from Advance micro Inc. and as the question specifically states that the purchase of additional 5,000 units will not lead to cancellation of order of part 88-461 to industrial division by admiral electric so the same was also assumed in case of Advance micro inc.

2) OPTION 1

If the industrial division’s sales to Admiral will be canceled if the commercial division does not buy from Admiral electric then the minimum transfer price charged by industrial division should be:

    Marginal cost = 7,75,000 (155 * 5,000 units)
Add: Contribution to be lost from reduced external sales = 2,50,000 (50*5000 units)
Contribution to be lost from cancellation of order of part 88-461 = 19,500 (WN 1)
     Minimum transfer price = 10,44,500
total units = 5,000
minimum transfer price per unit = 10,44,500 / 5,000 = $208.90

Total cost of additional 5,000 units = 5000 units * $208.90 / unit = $10,44,500

The desired change in transfer price = $208.90 - $205 = $3.90 per unit

WN 1: Sales value of part 88-461 = 95 per unit

Less: variable manufacturing cost = 65 per unit

Contribution = 30 per unit

total units sold = 650 unit

total contribution lost = 650 units*30 per unit = $19,500

OPTION 2

Purchase from Advance micro Inc. = $200 per unit

Total cost of additional 5,000 units = 5000 units * 200/ unit = $10,00,000

OPTION 3

Purchase from Admiral electric = $210 per unit.

Total cost of additional 5,000 units = 5000 units * 210 / unit = $10,50,000

In this case also it is advisable to purchase additional 5000 units from Advance micro Inc. as the total cost is least in this option.